A student earned grades of Upper A​, Upper F​, Upper A​, Upper D​, and Upper C. Those courses had the corresponding numbers of credit hours 4​, 4​, 2​, 4​, and 3. The grading system assigns quality points to letter grades as​ follows: Aequals​4; Bequals​3; Cequals​2; Dequals​1; Fequals0. Compute the grade point average​ (GPA) as a weighted mean and round the result with two decimal places. If the​ Dean's list requires a GPA of 3.00 or​ greater, did this student make the​ Dean's list? The grade point average is nothing. ​(Round to two decimal places as​ needed.) Did this student make the​ Dean's list? A. No because the students GPA is not 4.0 B. No because the​ student's GPA is less than 3.0 C. No because the student has at least one grade less than 3 D. Yes because at least two of the student grades are B or above
Answer #1

Total credit = 4 + 4 + 2 + 4 + 3 =17

Grade point average = (4 * 4 + 0 * 4 + 4 * 2 + 1 * 4 + 2 * 3) / 17 = 2

Option-B) No because the​ student's GPA is less than 3.0